# AniMaker: Automated Multi-Agent Animated Storytelling with MCTS-Driven Clip Generation

### v1

@@ Abstract.
Despite rapid advancements in video generation models, generating coherent, long-form storytelling videos that span multiple scenes and characters remains challenging. Current methods often rigidly convert pre-generated keyframes into fixed-length clips, resulting in disjointed narratives and pacing issues. Furthermore, the inherent instability of video generation models means that even a single low-quality clip can significantly degrade the entire output animation’s logical coherence and visual continuity. To overcome these obstacles, we introduce AniMaker, a multi-agent framework enabling efficient multi-candidate clip generation and storytelling-aware clip selection, thus creating globally consistent and story-coherent animation solely from text input. The framework is structured around specialized agents, including the Director Agent for storyboard generation, the Photography Agent for video clip generation, the Reviewer Agent for evaluation, and the Post-Production Agent for editing and voiceover, collectively realizing multi-character, multi-scene animation. Central to AniMaker’s approach are two key technical components: MCTS-Gen in Photography Agent, an efficient Monte Carlo Tree Search (MCTS)-inspired strategy that intelligently navigates the candidate space to generate high-potential clips while optimizing resource usage; and AniEval in Reviewer Agent, the first framework specifically designed for multi-shot animation evaluation, which assesses critical aspects such as story-level consistency, action completion, and animation-specific features by considering each clip in the context of its preceding and succeeding clips. Experiments demonstrate that AniMaker achieves superior quality as measured by popular metrics including VBench and our proposed AniEval framework, while significantly improving the efficiency of multi-candidate generation, pushing AI-generated storytelling animation closer to production standards.
@@
copyright: noneconference: ; ;
@@ 1. Introduction
The emergence of large-scale multimodal models (Yang et al., 2024; Kong et al., 2024; Liu et al., 2024c; Li et al., 2025) has ignited significant interest in storytelling videos. In response, a new wave of methods (Kim et al., 2024; Wu et al., 2024; Dalal et al., 2025; Guo et al., 2025) including MINT Video and TTT-Video has emerged, aiming to generate the entire video in a single pass. However, these methods still face significant challenges when generating long videos spanning multiple scenes and characters, particularly in maintaining visual continuity, narrative coherence, and non-repetitive content. In contrast, recent advancements in single-clip video generation, exemplified by models like Wan (Wan et al., 2025), Vidu(Bao et al., 2024), and Pika (Pika Labs, 2025), have pushed generated video clips closer to cinematic quality. Due to shorter clip duration, these models utilize textual and visual prompts directly and efficiently, improving motion quality, semantic fidelity, and consistency with reference images. Consequently, utilizing these models to generate a storytelling video with multiple clips appears highly promising.
Existing storytelling video frameworks (Lin et al., 2023; He et al., 2024; Li et al., 2024; Xu et al., 2025; Wu et al., 2025; Xie et al., 2024), which implement the composition of multiple video clips, generally follow a fixed pipeline: script → scene images → video clips → final video composition. While this modular format can effectively generate multi-shot long videos, it also introduces several limitations: Firstly, existing methods typically map scene images to fixed-length clips. It forms a rigid and fragmented video construction, leading to disjointed transitions and unnatural pacing, severely hindering the expressive continuity crucial for complex or extended actions. Secondly, due to the inherent instability of video generation models, a single flawed clip can noticeably degrade the overall video quality.
@@
To address these challenges, we introduce AniMaker, a multi-agent framework with MCTS-driven clip generation. This multi-agent framework mirrors professional production, including the Director Agent for storyboards, Photography Agent for clip generation, Reviewer Agent for evaluating the quality of clip candidates, and Post-Production Agent for editing the entire videos. These agents collaborate to enable automated multi-character, multi-scene storytelling without manual pre- or post-processing. The core Photography and Reviewer agents interact with each other under an MCTS-Gen scheme during clip generation. Specifically, based on Monte Carlo Tree Search (MCTS), MCTS-Gen offers an efficient strategy for navigating the vast candidate space of video generation. It strikes a balance between broad exploration and computational efficiency by intelligently allocating more generation opportunities to promising clips while encouraging the exploration of unexplored regions. For Reviewer Agent, we introduce a comprehensive evaluation framework AniEval, specifically designed for multi-shot storytelling animation. AniEval advances beyond metrics like VBench by implementing retrospective evaluation with cross-clip contextual references. It evaluates critical dimensions—story consistency, action completion, and animation-specific attributes—by analyzing each clip in the context of its preceding and succeeding clips. This contextual assessment serves as the quality evaluation mechanism for video-clip nodes within the MCTS-Gen framework.
We conduct extensive experiments on the dataset constructed from TinyStories (Eldan and Li, 2023), featuring complex interaction with multiple characters across diverse backgrounds. Experimental results demonstrate AniMaker’s superior performance across VBench, AniEval, and human evaluation, with significantly improved multi-candidate generation efficiency. This validates AniMaker’s effectiveness in bringing AI-generated storytelling videos closer to production-grade quality (as shown in Figure 1).
In summary, our main contributions are in three key aspects:
@@
  • Experimental results demonstrate AniMaker’s superior performance across all evaluation frameworks—achieving best scores across all scene image metrics, ranking 1st in VBench, 14.6% higher scores in our AniEval, and better human ratings (3.22 versus 2.07).
@@ 2. Related Work
Driven by the success of diffusion models, VDM (Ho et al., 2022) pioneers their application to video generation, leading to improved models like Stable Video Diffusion (Blattmann et al., 2023), ModelScope (Wang et al., 2023), VideoCrafter1 (Chen et al., 2023), and VideoCrafter2 (Chen et al., 2024). While these models excel at animating single text or image, they struggle with maintaining consistency across multiple video clips. Following Sora(Liu et al., 2024c), research has been increasingly focusing on coherent storytelling videos, with several works(Wu et al., 2024; Kim et al., 2024; Dalal et al., 2025; Guo et al., 2025) enhancing temporal modeling and contextual understanding. Despite these advancements, challenges persist as single-pass long video generation often suffers from inconsistencies and repetitive content over time. Considering this, models like HunyuanVideo (Kong et al., 2024), Wan (Wan et al., 2025), and CogVideoX (Yang et al., 2024) still prioritize the effective utilization of text prompts and reference images, resulting in more faithful generation and improved motion depiction. Industry solutions (Bao et al., 2024; Pika Labs, 2025; Pixverse AI, 2025; Runway AI, Inc., 2025) extend these advancements, enabling features like start/end frame-specific control.
@@
Storytelling video generation typically follows a pipeline: script → scene images → video clips → final video composition. Large multimodal models (LMMs) often act as high-level planners, coordinating each stage. Models such as VideoDirectorGPT (Lin et al., 2023), DreamStory (He et al., 2024), AnimDirector (Li et al., 2024), MM-StoryAgent (Xu et al., 2025), MovieAgent (Wu et al., 2025), and DreamFactory (Xie et al., 2024) adhere to this conventional pipeline. Notably, certain tasks like scriptwriting, character design, voiceovers, and sound effects are typically handled manually during pre- or post-processing.
@@ 3. Methodology
Our proposed framework, AniMaker, automates the creation of storytelling animation from text. It uses a multi-agent system that mirrors professional animation pipelines, incorporating novel components for efficient candidate generation and comprehensive evaluation. The overall architecture is detailed in Figure 2.
@@ 3.2. Pipeline Overview
AniMaker transforms textual input ($\mathbf{T}_{\text{prompt}}$) into compelling storytelling animation ($\mathbf{V}_{\text{final}}$) through four specialized agents working collaboratively. The Director Agent (Section 3.3.1) creates a detailed storyboard ($\boldsymbol{\mathcal{S}}_{\text{board}}$), the Photography Agent (Section 3.3.2) generates candidate video clips ($\mathbf{v}_{k}$) using MCTS-Gen (efficiently exploring generation space conserving computational resources), the Reviewer Agent (Section 3.3.3) evaluates clips with AniEval (our context-aware evaluation framework), and the Post-production Agent (Section 3.3.4) assembles clips with voiceover.
@@ 3.3.2. Photography Agent
Converting storyboards to clips in multi-shot AI video generation presents challenges including distorted appearances, inconsistent motion, and object inconsistencies. Inspired by filmmaking’s ”NG” concept and ”gacha” methods prevalent in AIGC, we recognize the necessity of generating multiple candidate clips to select the optimal one. For optimal selection, each clip must not only possess high individual quality but also ensure consistency and coherence with both preceding and succeeding clips. However, naively generating $k$ candidates per clip creates a combinatorial explosion (e.g., $k^{2}$ for two-clip sequences). Notably, poor-quality current clips allow for pruning the search space, as further exploration down such paths is unlikely to be satisfactory.
@@ 3.3.3. Reviewer Agent
Existing objective metrics, such as CLIP Score and Inception Score, may identify superior video generation models, but often struggle to differentiate among candidates from the same model using the same prompt. Similarly, the widely adopted VBench has significant limitations. For instance, some of its metrics, like ”dynamic degree,” are overly simplistic, merely measuring pixel changes rather than accurately reflecting character action. More critically, VBench’s ”Consistency” metrics, based on single-clip segmentation, prove unsuitable for multi-shot animation, which inherently involves frequent character and scene changes.
@@
In conclusion, AniEval comprises 4 primary domains with 14 fine-grained metrics for comprehensive assessment (Table 1). Additionally, AniEval supports contextual scoring by evaluating clips based on preceding and succeeding content, providing robust evaluation for multi-shot animation generation.
Table 1. AniEval Metrics.
@@ 3.3.4. Post-production Agent
The Post-production Agent transforms video clip sequences into a polished animation film through three stages. Firstly, it leverages Gemini 2.0 Flash to generate a detailed voiceover script specifying narration, dialogue, emotional tones, and desired voice timbres. The agent then selects appropriate voice profiles based on character attributes (age, gender) and assesses text length for audio-visual synchronization. Secondly, the script is processed through CosyVoice2 (Du et al., 2024) to generate audio tracks, which undergo verification for duration consistency and content accuracy. Finally, the agent employs the MoviePy library for film assembly, integrating validated subtitles and performing comprehensive editing to ensure precise synchronization between visuals, voiceover, and subtitles.
@@ 4.1.2. Baseline
Table 2. Scene Image evaluation on Contextual Coherence (Coherence), Image-Image Similarity (I-I Sim), and Text-Image Similarity (T-I Sim).
@@ 4.1.3. Evaluation Metrics
For scene image generation, we evaluate text-to-image alignment and cross-image consistency. Metrics include Text-to-Image CLIP (Coherence) (Radford et al., 2021), Image-to-Image Similarity (I-I Sim) (Gal et al., 2022), and Text-Image Similarity (T-I Sim) (Hessel et al., 2021). Video generation is assessed using VBench and our AniEval for comprehensive evaluation.
@@ 4.2. Qualitative Analysis
Based on samples in Figures 5 and 6, we qualitatively analyze AniMaker’s output, focusing on visual fidelity and narrative coherence in multi-character, multi-scene animated storytellings.
@@
AniMaker achieves superior visual consistency across scenes. While baselines struggle with character and background consistency during transitions, AniMaker maintains visual characteristics even when switching between different scenes. This stems from the Director Agent’s storyboard creation, providing consistent reference images for characters and backgrounds. Additionally, AniEval ensures coherent clip selection by evaluating cross-clip consistency metrics.
@@
AniMaker delivers smoother video transitions through effective generation and selection mechanisms. The Reviewer Agent uses AniEval, integrating cross-clip consistency metrics like DreamSim and applying contextual scoring across adjacent clips, minimizing visual disruptions.
@@
Table 3. VBench evaluation results, presenting scores for Image Quality (I.Q.), Semantic Consistency (S.C.), Background Consistency (B.C.), Animation Quality (A.Q.), Motion Smoothness (M.S.), Dynamic Degree (D.D.), and Average Rank (Rk. Avg. - the average ranking position across all models).
Model I.Q.$\uparrow$ S.C.$\uparrow$ B.C.$\uparrow$ A.Q.$\uparrow$ M.S.$\uparrow$ D.D.$\uparrow$ Rk. Avg.$\downarrow$
@@ 4.3. Quantitative Comparisons
Scene Image Generation Analysis
Scene images (Keyframes) are the core output of a storyboard. Table 2 demonstrates that AniMaker outperforms all competing methods across metrics. Notably, AniMaker achieves a Text-to-Image Similarity (T-I Sim) score of 0.31, representing a 19.2% improvement over the best-performing baseline. This advantage stems from our multimodal storyboard generation approach that incorporates character references, background references, and text prompts for keyframe generation, rather than relying solely on text.
@@
We introduce AniEval for better evaluation of storytelling Animation. Table 4 presents results from our AniEval framework. AniMaker outperforms all competitors with a total score of 76.72, representing a 14.6% improvement over the second-best model (VideoGen-of-Thought at 66.93). Particularly noteworthy is AniMaker’s exceptional performance in Video Consistency (V.C.), surpassing the best-performing baseline by 15.5%. AniMaker’s relatively lower performance in Text-Video Alignment (T.V.A.) can be attributed to our agent framework’s creative adaptation of stories into scripts, where additional narrative elements are introduced. Compared to VBench, AniEval’s assessment results align remarkably better with human evaluations (Table 5), confirming that AniEval provides a more comprehensive and accurate assessment of multi-shot animation quality than previous metrics.

### v2

# AniMaker: Multi-Agent Animated Storytelling with MCTS-Driven Clip Generation

@@ Abstract.
Despite rapid advancements in video generation models, generating coherent, long-form storytelling videos that span multiple scenes and characters remains challenging. Current methods often rigidly convert pre-generated keyframes into fixed-length clips, resulting in disjointed narratives and pacing issues. Furthermore, the inherent instability of video generation models means that even a single low-quality clip can significantly degrade the entire output animation’s logical coherence and visual continuity. To overcome these obstacles, we introduce AniMaker, a multi-agent framework enabling efficient multi-candidate clip generation and storytelling-aware clip selection, thus creating globally consistent and story-coherent animation solely from text input. The framework is structured around specialized agents, including the Director Agent for storyboard generation, the Photography Agent for video clip generation, the Reviewer Agent for evaluation, and the Post-Production Agent for editing and voiceover, collectively realizing multi-character, multi-scene animation. Central to AniMaker’s approach are two key technical components: MCTS-Gen in Photography Agent, an efficient Monte Carlo Tree Search (MCTS)-inspired strategy that intelligently navigates the candidate space to generate high-potential clips while optimizing resource usage; and AniEval in Reviewer Agent, the first framework specifically designed for multi-shot animation evaluation, which assesses critical aspects such as story-level consistency, action completion, and animation-specific features by considering each clip in the context of its preceding and succeeding clips. Experiments demonstrate that AniMaker achieves superior quality as measured by popular metrics including VBench and our proposed AniEval framework, while significantly improving the efficiency of multi-candidate generation, pushing AI-generated storytelling animation closer to production standards. Code and data for this paper are at https://animaker-dev.github.io/
@@
journalyear: 2025copyright: acmlicensedconference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3764009isbn: 979-8-4007-2137-3/2025/12ccs: Applied computing Arts and humanitiesccs: Computing methodologies Computer visionccs: Computing methodologies Animation
@@ 1. Introduction
The emergence of large-scale video generation models (Yang et al., 2024; Kong et al., 2024; Liu et al., 2024c) has ignited significant interest in storytelling videos. In response, a new wave of methods (Kim et al., 2024; Wu et al., 2024; Dalal et al., 2025; Guo et al., 2025; Xu et al., 2025b) including MINT Video and TTT-Video has emerged, aiming to generate the entire video in a single pass. However, these methods still face significant challenges when generating long videos spanning multiple scenes and characters, particularly in maintaining visual continuity, narrative coherence, and non-repetitive content. In contrast, recent advancements in single-clip video generation, exemplified by models like Vidu(Bao et al., 2024), Wan (Wan et al., 2025), and Pika (Pika Labs, 2025), have pushed generated video clips closer to cinematic quality. Due to shorter clip duration, these models utilize textual and visual prompts directly and efficiently, improving motion quality, semantic fidelity, and consistency with reference images. Consequently, utilizing these models to generate a storytelling video with multiple clips appears highly promising.
Existing storytelling video frameworks (Lin et al., 2023; He et al., 2024; Xie et al., 2024; Li et al., 2024; Xu et al., 2025a; Wu et al., 2025), which implement the composition of multiple video clips, generally follow a fixed pipeline: script → keyframes → video clips → final video composition. While this modular format can effectively generate multi-shot long videos, it also introduces several limitations: Firstly, existing methods typically map keyframes to fixed-length clips. It forms a rigid and fragmented video construction, leading to disjointed transitions and unnatural pacing, severely hindering the expressive continuity crucial for complex or extended actions (Jiang et al., 2021). Secondly, due to the inherent instability of video generation models, a single flawed clip can noticeably degrade the overall video quality.
@@
To address these challenges, we introduce AniMaker, a multi-agent framework with MCTS-driven clip generation. This framework mirrors professional production (Xu et al., 2025b), including the Director Agent for storyboards construction, Photography Agent for video clip generation, Reviewer Agent for evaluating the quality of video clip candidates, and Post-Production Agent for editing and assembling the entire sequence of videos. These agents collaborate to enable automated multi-character, multi-scene storytelling without manual pre- or post-processing. The core Photography and Reviewer agents interact under a MCTS-Gen scheme during clip generation. Specifically, based on Monte Carlo Tree Search (MCTS), MCTS-Gen offers an efficient strategy for navigating the vast candidate space of video generation. It strikes a balance between broad exploration and computational efficiency by intelligently allocating more generation opportunities to promising clips while encouraging the exploration of unexplored regions. For Reviewer Agent, we introduce a comprehensive evaluation framework AniEval, specifically designed for multi-shot storytelling (Cailin Zhuang, 2025) animation. AniEval advances beyond metrics like VBench by implementing retrospective evaluation with cross-clip contextual references. It evaluates critical dimensions—story consistency, action completion, and animation-specific attributes—by analyzing each clip in the context of its preceding and succeeding clips. This contextual assessment serves as the quality evaluation mechanism for video-clip nodes within the MCTS-Gen framework.
We conduct extensive experiments on the dataset constructed from TinyStories (Eldan and Li, 2023), featuring complex interaction with multiple characters across diverse backgrounds. Experimental results demonstrate AniMaker’s superior performance across VBench, AniEval, and human evaluation, with significantly improved multi-candidate generation efficiency. This validates AniMaker’s effectiveness in bringing AI-generated storytelling videos closer to production-grade quality (as shown in Figure 1). With our current work successfully targeting simple storytelling, we believe further development is essential to produce more complex and sophisticated animation styles.
In summary, our main contributions are:
@@
  • We experimentally demonstrate Animaker’s superior performance across all evaluation frameworks—achieving best scores across all keyframe evaluation metrics, ranking 1st in VBench, 14.6% higher scores in our AniEval, and better human ratings (3.22 versus 2.07).
@@ 2. Related Work
Driven by the success of diffusion models, VDM (Ho et al., 2022) pioneers their application to video generation, leading to improved models like Stable Video Diffusion (Blattmann et al., 2023), ModelScope (Wang et al., 2023b), VideoCrafter1 (Chen et al., 2023), and VideoCrafter2 (Chen et al., 2024). While these models excel at animating single text or image, they struggle with maintaining consistency across multiple video clips. The emergence of Sora (Liu et al., 2024c) has spurred a demand for coherent, consistent, and controllable storytelling videos. In response, several works(Wu et al., 2024; Kim et al., 2024; Dalal et al., 2025; Guo et al., 2025) have emerged to enhance storytelling video generation by improving temporal modeling and contextual understanding. Despite these advancements, challenges persist as single-pass long video generation often suffers from inconsistencies and repetitive content over time. Considering this, models like HunyuanVideo (Kong et al., 2024), Wan (Wan et al., 2025), and CogVideoX (Yang et al., 2024) still prioritize the effective utilization of text prompts and reference images, resulting in more faithful generation and improved motion depiction. Industry solutions (Bao et al., 2024; Pika Labs, 2025; Pixverse AI, 2025; Runway AI, Inc., 2025) extend these advancements, enabling features like start/end frame-specific control.
@@
Storytelling video generation typically follows a pipeline: script → keyframes → video clips → final video composition. Large multimodal models (LMMs) often act as high-level planners to coordinate each stage (e.g., (Google Cloud, 2025; Hurst et al., 2024; Li et al., 2025)). Frameworks such as VideoDirectorGPT (Lin et al., 2023), DreamStory (He et al., 2024), AnimDirector (Li et al., 2024), MM-StoryAgent (Xu et al., 2025a), MovieAgent (Wu et al., 2025), and DreamFactory (Xie et al., 2024) adhere to this conventional pipeline. Notably, certain tasks like scriptwriting, character design, voiceovers, and sound effects are typically handled manually during pre- or post-processing.
@@ 3. Methodology
Our proposed framework, AniMaker, automates the creation of foundational storytelling animation from text. It uses a multi-agent system that mirrors professional animation pipelines, incorporating novel components for efficient candidate generation and comprehensive evaluation. The overall architecture is detailed in Figure 2.
@@ 3.2. Pipeline Overview
AniMaker transforms textual input ($\mathbf{T}_{\text{prompt}}$) into compelling storytelling animation ($\mathbf{V}_{\text{final}}$) through four specialized agents working collaboratively. Central to this pipeline is the use of LLM prompt engineering by its four specialized agents, and we refer to the appendix for related details. The Director Agent (Section 3.3.1) creates a detailed storyboard ($\boldsymbol{\mathcal{S}}_{\text{board}}$), the Photography Agent (Section 3.3.2) generates candidate video clips ($\mathbf{v}_{k}$) using MCTS-Gen (efficiently exploring generation space conserving computational resources), the Reviewer Agent (Section 3.3.3) evaluates clips with AniEval (our context-aware evaluation framework), and the Post-production Agent (Section 3.3.4) assembles clips with voiceovers and subtitles.
@@ 3.3.2. Photography Agent
Converting storyboards to clips in multi-shot AI video generation presents challenges including distorted appearances, inconsistent motion, and object inconsistencies. Drawing from the ”No Good” (NG) process in filmmaking, where numerous takes are recorded to achieve the perfect shot, we recognize the need to produce multiple candidate clips to identify the optimal one. For optimal selection, each clip must not only possess high individual quality but also ensure consistency and coherence with both preceding and succeeding clips. However, naively generating $k$ candidates per clip creates a combinatorial explosion (e.g., $k^{2}$ for two-clip sequences). Notably, poor-quality current clips allow for pruning the search space, as further exploration down such paths is unlikely to be satisfactory.
@@ 3.3.3. Reviewer Agent
Existing objective metrics, such as CLIP Score and Inception Score, may identify superior video generation models, but often struggle to differentiate among candidates generated by the same model. Similarly, the widely adopted VBench has significant limitations. For instance, some of its metrics, like ”dynamic degree”, are overly simplistic, merely measuring pixel changes rather than accurately reflecting character action. More critically, VBench’s ”Consistency” related metrics, based on single-clip segmentation, prove unsuitable for multi-shot animation, which inherently involves frequent character and scene changes.
@@
In conclusion, AniEval comprises 4 primary domains with 14 fine-grained metrics for comprehensive assessment (Table 1). Additionally, AniEval supports contextual scoring by evaluating clips based on preceding and succeeding content, providing robust evaluation for multi-shot animation generation. Further implementation details of AniEval can be found in the appendix.
Table 1. Introduction of AniEval Metrics.
@@ 3.3.4. Post-production Agent
The Post-production Agent transforms video clip sequences into a polished animation film through three stages. Firstly, it leverages Gemini 2.0 Flash to generate a detailed voiceover script specifying narration, dialogue, emotional tones, and desired voice timbres. The agent then selects appropriate voice profiles based on character attributes (age, gender) and assesses text length for audio-visual synchronization. Secondly, the script is processed through CosyVoice2 (Du et al., 2024) to generate audio tracks, which undergo verification regarding duration consistency and content accuracy. Finally, the agent employs the MoviePy library for film assembly, integrating validated subtitles and performing comprehensive editing to ensure precise synchronization between visuals, voiceovers, and subtitles.
@@ 4.1.2. Baseline
Table 2. Keyframe evaluation on Contextual Coherence (Coherence), Image-Image Similarity (I-I Sim), and Text-Image Similarity (T-I Sim).
@@ 4.1.3. Evaluation Metrics
For keyframe generation, we evaluate text-to-image alignment and cross-image consistency. Metrics include Text-to-Image CLIP (Coherence) (Radford et al., 2021a), Image-to-Image Similarity (I-I Sim) (Gal et al., 2022), and Text-Image Similarity (T-I Sim) (Hessel et al., 2021). Video generation is assessed using VBench and our AniEval for comprehensive evaluation.
@@ 4.2. Qualitative Analysis
Based on the samples depicted in Figures 5 and 6, we conduct a qualitative analysis of AniMaker’s output, with a focus on visual fidelity and narrative coherence in multi-character, multi-scene animated storytelling.
@@
AniMaker demonstrates exceptional visual consistency across scenes. While baseline methods often face challenges in maintaining character and background consistency during scene transitions, AniMaker successfully preserves visual characteristics, even when switching back and forth between distinct scenes. This can be attributed to the Director Agent’s storyboard creation process, which provides a reliable set of reference images for both characters and backgrounds. Additionally, AniEval plays a crucial role in ensuring narrative coherence by incorporating a meticulous clip selection process, which evaluates consistency related metrics across clips.
@@
AniMaker achieves smoother video transitions through an effective generation and selection mechanism. Central to this process is the Reviewer Agent, which leverages our AniEval framework to maximize visual continuity. By integrating cross-clip consistency metrics such as DreamSim and applying contextual scoring to adjacent clips, the agent effectively minimizes jarring visual disruptions between shots.
@@
Table 3. VBench evaluation results, presenting scores for Image Quality (I.Q.), Semantic Consistency (S.C.), Background Consistency (B.C.), Animation Quality (A.Q.), Motion Smoothness (M.S.), Dynamic Degree ($D.$), and Average Rank (Rk. Avg. - the average ranking position across all models).
[…]
@@ 4.3. Quantitative Comparisons
Keyframe Generation Analysis
Keyframes are the core components of a storyboard. Table 2 demonstrates that AniMaker outperforms all other competing methods across metrics. Notably, AniMaker achieves a Text-to-Image Similarity (T-I Sim) score of 0.31, representing a 19.2% improvement over the best-performing baseline method. This advantage stems from our multimodal storyboard generation approach that incorporates character references, background references, and text prompts for keyframe generation, rather than relying solely on text.
@@
We introduce AniEval for better evaluation of storytelling Animation. Table 4 presents results from our AniEval framework. AniMaker outperforms all competitors with a total score of 76.72, representing a 14.6% improvement over the second-best method (i.e. VideoGen-of-Thought with 66.93). Particularly noteworthy is AniMaker’s exceptional performance in Video Consistency (V.C.), surpassing the best-performing baseline method by 15.5%. While, AniMaker’s relatively lower performance in Text-Video Alignment (T.V.A.) can be attributed to our agent framework’s creative adaptation of stories into scripts, where additional narrative elements are introduced. Compared to VBench, AniEval’s assessment results align remarkably better with human evaluations (Table 5), demonstrating that AniEval provides a more comprehensive and accurate assessment of multi-shot animation quality than previous metrics.